OpenStudy (anonymous):

choose the new equation after the expression equivalent to x from the second equation is substituted into the first equation of these two3x + 2y = –21and x – 3y = 4. Here are the choices. A)3x + 2(3y + 4) = –21 B)3x + 2(–3y + 4) = –21 C)3(3y + 4) + 2y = –21 D)3(–3y + 4) + 2y = –21 I need a quick answer please and thank you :)

OpenStudy (anonymous):

C) 3(3y + 4) + 2y = –21 is the right option

OpenStudy (anonymous):

3x + 2y = –21and x – 3y = 4 ==>x = 3y + 4 so first becomes (on substitution) 3(3y + 4) + 2y = –21

OpenStudy (anonymous):

Thanks alot for the help :)

OpenStudy (anonymous):

welcome ....☺

hero (hero):

Needhelp, I responded to the previous question you posted.

OpenStudy (anonymous):

oh thank you too :)